OpenStudy (anikate):

a bag contains 2 green marbles, 4 blue marbles, and 5 red marbles. What is the probability of choosing 2 blue marbles, if the first marble is returned to the bag before the second marble is chosen?

OpenStudy (anonymous):

total number of marbles is 11 and the probability that the first one is blue is \[\frac{4}{11}\] also the probability that the second one is blue is \[\frac{4}{11}\] and you want the probability that they are both blue. answer is \[\frac{4}{11}\times \frac{4}{11}\]

OpenStudy (anonymous):

make a tree diagram it's way easier to understand what's happening here.

OpenStudy (anonymous):

if the first one was not returned to the bag it would be \[\frac{4}{11}\times \frac{3}{10}\]

OpenStudy (anikate):

slow down i dont understand anything

OpenStudy (anonymous):

i will help you anikate

OpenStudy (anikate):

k thx

OpenStudy (anonymous):

you can make a tree, but it is a fact that since the events are independent (meaning the outcome of one doesn't change the outcome of the other) that \[P(A\cap B)=P(A)\times P(B)\]

OpenStudy (anonymous):

memorizing formulas won't get you anywhere in life lol

OpenStudy (anikate):

i am just 7th grade

OpenStudy (anonymous):

true but if you want to compute things you will need to know under what circumstances to add, subtract, multiply or divide

OpenStudy (anikate):

plz help me with my question

OpenStudy (anonymous):

how many marbles in the bag total?

OpenStudy (anonymous):

start out with 3 lines...one for green one for blue and one for red. mark on each line the probability for that specific color on the first pick. for example, red would be 5/11, blue 4/11, green 2/11.

OpenStudy (anonymous):

since you are replacing the marble you took, the probability doesn't change on the second pick

OpenStudy (anonymous):

2

OpenStudy (anikate):

hello?

OpenStudy (anikate):

plz help me

OpenStudy (anonymous):

anikate i answered you, look up

OpenStudy (anikate):

nothing came up

OpenStudy (anikate):

hello?

OpenStudy (anonymous):

start out with 3 lines...one for green one for blue and one for red. mark on each line the probability for that specific color on the first pick. for example, red would be 5/11, blue 4/11, green 2/11. since you are replacing the marble you took, the probability doesn't change on the second pick. it's 4/11X4/11. So 16/121

OpenStudy (anikate):

still got nothing

OpenStudy (anonymous):

Ok, look. You don't need to worry about the probabilities for the red and green marbles since it's only the blues you care about. Just compute the probability of choosing a blue marble first.

OpenStudy (anonymous):

How many blue marbles are in the bag?

OpenStudy (anikate):

2

OpenStudy (anonymous):

"a bag contains 2 green marbles, 4 blue marbles, and 5 red marbles." There are 4 blue marbles right?

OpenStudy (anikate):

no 2 blue

OpenStudy (anonymous):

Type the problem again?

OpenStudy (anikate):

a bag contains 2 green marbles, 4 blue marbles, and 5 red marbles. What is the probability of choosing 2 blue marbles if the first marble is returned to the bag before the second marble is chosen?

OpenStudy (anikate):

um hello?

OpenStudy (anonymous):

ready?

OpenStudy (anikate):

yes

OpenStudy (anonymous):

we do this in 3 easy steps

OpenStudy (anonymous):

how many marbles in the bag total?

OpenStudy (anikate):

11 marbles

OpenStudy (anonymous):

right. how many are blue?

OpenStudy (anikate):

4 blue marble

OpenStudy (anonymous):

exactly. now what is the probability that if you pick one at random you get a blue one?

OpenStudy (anikate):

4/11

OpenStudy (anonymous):

which is the same question as "what is the ratio of the number of blue marbles to the total number of marbles?" yes you have it!

OpenStudy (anikate):

but

OpenStudy (anonymous):

so the probability the first marble is 4/11

OpenStudy (anikate):

o

OpenStudy (anonymous):

we are not done yet, but almost

OpenStudy (anikate):

wait

OpenStudy (anonymous):

i wait

OpenStudy (anikate):

4/ll for blue

OpenStudy (anikate):

right

OpenStudy (anonymous):

yes

OpenStudy (anikate):

k

OpenStudy (anonymous):

site is slow be patient

OpenStudy (anonymous):

ok now if you are going to pick two blue marbles, both the first and second have to be blue. that is clear yes?

OpenStudy (anikate):

yup

OpenStudy (anonymous):

so you know your first marble was blue, otherwise you cannot pick 2 blue ones yes?

OpenStudy (anonymous):

now you have picked a blue marble. you put it back in the bag. how many marbles are in the bag?

OpenStudy (anikate):

11

OpenStudy (anonymous):

how many are blue?

OpenStudy (anikate):

4

OpenStudy (anonymous):

yes what is the probability you pick a blue marble on the second try?

OpenStudy (anikate):

4/11

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

ok now we finish the question. in order for them BOTH to be blue we must pick a blue marble on the first try AND pick a blue marble on the second try yes?

OpenStudy (anonymous):

Good stuff satellite.

OpenStudy (anonymous):

yeah well now comes the tricky part!

OpenStudy (anonymous):

for the probability that BOTH are blue we have to take the probability of each and multiply them together!

OpenStudy (anikate):

16/121

OpenStudy (anonymous):

our last job is to take \[\frac{4}{11}\times \frac{4}{11}\]

OpenStudy (anonymous):

that is what i get, yes. you have it and we are done unless you want to do another one

OpenStudy (anikate):

i have other questions

OpenStudy (anonymous):

ask one here and we can do it together

OpenStudy (anikate):

k

OpenStudy (anikate):

here it goes

OpenStudy (anikate):

there is a prisoner

OpenStudy (anikate):

and there are 3 routes for her to escape

OpenStudy (anikate):

route A has 1/4 chances of being successful

OpenStudy (anikate):

route b is 1/5

OpenStudy (anikate):

route c is 1/6

OpenStudy (anonymous):

got it so far

OpenStudy (anikate):

and

OpenStudy (anikate):

the probability of selecting each route is 1/3

OpenStudy (anikate):

what is the probability that she might successfully escape by using road B

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

we proceed just like the last one.

OpenStudy (anikate):

k

OpenStudy (anonymous):

first if she is going to get out using road B two things must happen. first she must choose road B yes?

OpenStudy (anikate):

yes

OpenStudy (anonymous):

what is that probability?

OpenStudy (anikate):

probability of selecting each route is 1/3

OpenStudy (anikate):

2 minutes i will be right back

OpenStudy (anonymous):

ok so it is 1/3 then if she select road B what is the probability that she escapes?

OpenStudy (anonymous):

ok write when you get back

OpenStudy (anikate):

1/5

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

now what is the probability that they BOTH occur?

OpenStudy (anikate):

huh?

OpenStudy (anonymous):

first the prisoner has to pick the road. then he has to escape. BOTH have to happen yes?

OpenStudy (anonymous):

same as previous problem where BOTH marbles have to be blue

OpenStudy (anikate):

yes